OpenStudy (anonymous):

What is the equation of the directrix of the parabola given by the equation y2 = -24x? y = 4 x = 4 y = 6 x = 6

OpenStudy (anonymous):

@dan815

OpenStudy (welshfella):

Hint:- for the parabola y^2 = 4ax the directrix is x = -a

OpenStudy (anonymous):

i have no idea how to do this

OpenStudy (welshfella):

compare the 2 functions y^2 = 4 a x directrix is x = -a y^2 = -24 x = 4 (-6) x directrix is x = -(-6) x = 6